# Time-harmonic waves in Korteweg and nematic-Korteweg fluids

### v1

@@ Abstract
$\underline{\underline{\sigma}}^{(V)}=p\underline{\underline{I}}-u_{1}\rho\left(\nabla\rho\otimes\nabla\rho\right)+u_{2}\left(\nabla\rho\cdot\vec{n}\right)\nabla\rho\otimes\vec{n},$ (3)
@@
Virga has given a full derivation of the stress tensor (3) [5].
A classical derivation of Korteweg’s model can be found in Toupin [6, 7], where continuum equations are derived from a classical principle of virtual work. The Korteweg equations can also be derived from a non-standard variational principle [8], can be cast within the framework of hyperelasticity [9], and derived from the kinetic theory of gases [10].
@@ I Helmholtz–Korteweg equation
$-\omega^{2}S(\vec{x})-c_{0}^{2}\Delta S(\vec{x})+\rho_{0}^{2}u_{1}\Delta^{2}S(\vec{x})=0.$ (12)
@@
$W(\rho,\nabla\rho)=c_{0}^{2}\rho+\frac{1}{2}u_{1}\lVert\nabla\rho\rVert^{2}.$ (13)
@@
    $c_{0}^{2}\rho_{0}S(\vec{x})-\rho_{0}^{3}u_{1}\Delta S(\vec{x})=0,$ (16)
    thus the sound-soft boundary conditions correspond to imposing homogeneous Dirichlet boundary conditions on $S(\vec{x})$ and on $\Delta S(\vec{x})$.
@@
    Sound-hard boundary conditions, which correspond to imposing that the normal derivative of the fluid velocity $\partial_{\vec{\nu}}\vec{v}$ vanishes along $\Gamma$. Using (7) and assuming the fluid velocity is also time-harmonic, we get
    $i\omega\rho_{0}(\vec{n}\cdot\vec{\nu})=c_{0}^{2}\partial_{\vec{\nu}}S(\vec{x})-\rho_{0}^{2}u_{1}\partial_{\vec{\nu}}\Delta S(\vec{x}),$ (17)
    thus the sound-hard boundary conditions correspond to imposing homogeneous Neumann boundary conditions on $S(\vec{x})$ and on $\Delta S(\vec{x})$.
@@
    Impedance boundary conditions, which correspond to imposing that the normal derivative of the fluid velocity $\partial_{\vec{\nu}}\vec{v}$ is proportional to the excess pressure along $\Gamma$. This boundary condition is equivalent to imposing Robin boundary conditions on $S(\vec{x})$ and on $\Delta S(\vec{x})$. The proportionality constant is usually denoted by $\zeta$ and is called the impedance of the boundary.
@@ II Nematic Helmholtz–Korteweg equation
where $c_{1},c_{2}>0$ are constitutive constants. This approach was extended to the case of a general acoustic wave in Selinger et al. [25], where the elastic energy augmentation is assumed to couple the nematic director field with the gradient of the density field. In this context Virga’s theory regards liquid crystals as Korteweg fluids and proposes the constitutive relation (3) for the Cauchy stress tensor, which in a hyperelastic framework can be derived from the elastic energy density
$W(\rho,\nabla\rho,\vec{n})=c_{0}^{2}\rho+\frac{1}{2}u_{1}\lVert\nabla\rho\rVert^{2}+\frac{1}{2}u_{2}(\nabla\rho\cdot\vec{n})^{2}.$ (19)
The acoustic energy $W_{a}$ is derived from $W$ by assuming that the density $\rho$ represents a planar wave.
@@
    Sound-hard boundary conditions also change since the normal derivative of the fluid velocity $\partial_{\vec{\nu}}\vec{v}$ now satisfies the equation
    $i\omega\rho_{0}(\vec{n}\cdot\vec{\nu})=c_{0}^{2}\partial_{\vec{\nu}}S(\vec{x})$ $-\rho_{0}^{2}u_{1}\partial_{\vec{\nu}}\Delta S(\vec{x})$ (30)
@@ III Plane waves
$-1+\kappa^{2}+\frac{1}{4}\tau_{1}^{2}\omega^{2}\kappa^{4}=0,\qquad\tau_{1}\coloneqq 4\frac{\rho_{0}\sqrt{u_{1}}}{c_{0}^{2}}.$ (37)
@@
$-1+\left(\frac{c_{0}}{c}\right)^{2}+\frac{1}{4}\tau_{1}^{2}\omega^{2}\left(\frac{c_{0}}{c}\right)^{4}=0.$ (43)
@@
$-1-\left(\frac{c_{0}}{c}\right)^{2}\alpha+\frac{1}{4}\tau_{1}^{2}\omega^{2}\left(\frac{c_{0}}{c}\right)^{4}\alpha^{2}=0.$ (46)
@@
$-\omega^{2}+c_{0}^{2}k^{2}+\rho_{0}^{2}u_{1}k^{4}+\rho_{0}^{2}u_{2}k^{4}\left(\vec{d}\cdot\vec{n}\right)^{2}=0,$ (50)
@@
$-\omega^{2}+c_{0}^{2}k^{2}+\left[\rho_{0}^{2}u_{1}+\rho_{0}^{2}u_{2}\cos^{2}(\xi)\right]k^{4}=0.$ (51)
@@
$-1+\kappa^{2}+\frac{1}{4}\tau_{2}^{2}\omega^{2}\kappa^{4}=0,$ (53)
@@ IV Plane wave reflection
    For sound-soft boundary conditions along the real axis we have $s_{0}e^{ikd_{1}x_{1}}+s_{0}Ae^{ikd_{1}x_{1}}=0$, thus we need to impose $A=-1$.
@@
Figure 2: An acoustic reflection phenomenon in a nematic Korteweg fluid can be caused by a discontinuity in the nematic director field. (Top) The nematic director field is parallel to the wave-vector on the left side of the line marked in green, and orthogonal on the right side. As a result, a plane wave is partially reflected along the line. (Bottom) No reflection occurs when the nematic director has no discontinuity.
@@ V Scattering by a circular obstacle
$-c_{0}^{2}\Delta S^{+}_{\ell}(\vec{\xi})+\Delta^{2}S^{+}_{\ell}(\vec{\xi})+\gamma^{-1}\nabla\!\cdot\!\nabla\left[{\vec{n}\cdot\underline{\underline{\mathcal{H}S^{+}_{\ell}}}\vec{n}}\right]=0.$ (66)
@@
$-c_{0}S^{+}_{\ell}(\vec{\xi})+\Delta S^{+}_{\ell}(\vec{\xi})+\gamma^{-1}{\vec{n}\cdot\underline{\underline{\mathcal{H}S^{+}_{\ell}}}\vec{n}}=0,$ (67)
@@
$-c_{0}S^{+}_{\ell}(\vec{\xi})+\nabla\cdot\left[(I+\gamma^{-1}\vec{n}\otimes\vec{n})\nabla S^{+}_{\ell}\right]=0.$ (68)
@@
where $(r,\theta)$ are the polar coordinates of $\vec{x}$, $R$ is the radius of the circular obstacle, $H_{j}^{(1)}$ are the Hankel functions of the first kind, $k$ is the wave number, and $a_{j}$ are the coefficients obtained by expanding the incoming plane wave $S^{-}$ in circular harmonics via the Jacobi–Anger formula [29, eq. (22)], i.e.
@@
where $J_{j}$ is the Bessel function of the first kind and $\psi$ is the angle of incidence of the incoming plane wave. Notice now that the wave number $k$ appearing in (74) depends on the incoming plane wave and needs to satisfy the dispersion relation (50). In particular, as discussed in Section III, the wave number $k$ is smaller when the wave propagates orthogonal to the nematic director field. In regions where the incident plane wave is orthogonal to the nematic director field we can expect the amplitude of the reflected wave to be smaller, given the following asymptotic expansion of $S^{+}$ [29, eq. (38)]:
@@ Acknowledgements.
This work was funded by the Engineering and Physical Sciences Research Council [grant numbers EP/R029423/1 and EP/W026163/1]. The authors would like to thank E. Virga for helpful discussions and comments on the first draft of this manuscript. The authors would like to thank also T. van Beeck and C. Lehrenfeld for their help in understanding the correct boundary conditions for the Helmholtz–Korteweg equation. For the purpose of open access, the author has applied a CC BY public copyright licence to any Author Accepted Manuscript (AAM) version arising from this submission. No new data were generated or analysed during this work.

### v2

@@ Abstract
$\underline{\underline{\sigma}}^{(V)}=p\underline{\underline{I}}-u_{1}\rho\left(\nabla\rho\otimes\nabla\rho\right)-u_{2}{\color[rgb]{0.2980392156862745,0.4470588235294118,0.6901960784313725}\definecolor[named]{pgfstrokecolor}{rgb}{0.2980392156862745,0.4470588235294118,0.6901960784313725}\rho}\left(\nabla\rho\cdot\vec{n}\right)\nabla\rho\otimes\vec{n},$ (3)
@@
While general second-grade materials can convey internal torques by a hyperstress, in Korteweg fluids this vanishes identically [5].
The state variables for a Korteweg fluid are its density $\rho$, its gradient $\nabla\rho$, and (in the context of nematic-Korteweg fluids) also the nematic director $\vec{n}$. Under the assumption that, given an increasing function $\sigma_{0}(\rho)$, the Helmholtz free energy density of the fluid is given by
$W(\rho,\nabla\rho,\vec{n})=\sigma_{0}(\rho)+\frac{1}{2}u_{1}\lVert\nabla\rho\rVert^{2}+\frac{1}{2}u_{2}(\nabla\rho\cdot\vec{n})^{2},$ (5)
the Korteweg and nematic-Korteweg stress tensors can be derived from a hyperelastic framework [6]. A more classical derivation of Korteweg’s model can be found in Toupin [7, 8], where continuum equations are derived from a classical principle of virtual work. The Korteweg equations can also be derived from a non-standard variational principle [9] and from the kinetic theory of gases [10].
@@ I Helmholtz–Korteweg equation
$\omega^{2}S(\vec{x})-c_{0}^{2}\Delta S(\vec{x})+\rho_{0}^{2}u_{1}\Delta^{2}S(\vec{x})=0.$ (13)
@@
$W(\rho,\nabla\rho)={\color[rgb]{0.2980392156862745,0.4470588235294118,0.6901960784313725}\definecolor[named]{pgfstrokecolor}{rgb}{0.2980392156862745,0.4470588235294118,0.6901960784313725}\frac{1}{2}c_{0}^{2}\rho^{2}}+\frac{1}{2}u_{1}\lVert\nabla\rho\rVert^{2}.$ (14)
@@
    $c_{0}^{2}\rho_{0}S(\vec{x})-\rho_{0}^{3}u_{1}\Delta S(\vec{x})=0.$ (17)
    Thus, assuming the sound-soft boundary conditions for the classical Helmholtz equations also apply, we obtain that the sound-soft boundary conditions correspond to imposing homogeneous Dirichlet boundary conditions on $S(\vec{x})$ and on $\Delta S(\vec{x})$.
@@
[…]
    ${\color[rgb]{0.2980392156862745,0.4470588235294118,0.6901960784313725}\definecolor[named]{pgfstrokecolor}{rgb}{0.2980392156862745,0.4470588235294118,0.6901960784313725}-i\omega\rho_{0}(\vec{V}\cdot\vec{\nu})}=c_{0}^{2}\partial_{\vec{\nu}}S(\vec{x})-\rho_{0}^{2}u_{1}\partial_{\vec{\nu}}\Delta S(\vec{x}).$ (18)
    Thus, assuming that sound-hard boundary conditions for the classical Helmholtz equation also apply, we obtain that the sound-hard boundary conditions correspond to imposing homogeneous Neumann boundary conditions on $S(\vec{x})$ and on $\Delta S(\vec{x})$.
@@
[…]
    $c_{0}^{2}\rho_{0}S(\vec{x})-\rho_{0}^{3}u_{1}\Delta S(\vec{x})$ $={\color[rgb]{0.2980392156862745,0.4470588235294118,0.6901960784313725}\definecolor[named]{pgfstrokecolor}{rgb}{0.2980392156862745,0.4470588235294118,0.6901960784313725}i\zeta}c_{0}^{2}\partial_{\vec{\nu}}S(\vec{x})$ (19)
    $-{\color[rgb]{0.2980392156862745,0.4470588235294118,0.6901960784313725}\definecolor[named]{pgfstrokecolor}{rgb}{0.2980392156862745,0.4470588235294118,0.6901960784313725}i\zeta}\rho_{0}^{2}u_{1}\partial_{\vec{\nu}}\Delta S(\vec{x}).$
    Assuming the impedance boundary conditions for the classical Helmholtz equation also apply, we observe that impedance boundary conditions are equivalent to imposing Robin boundary conditions on $S(\vec{x})$ and on $\Delta S(\vec{x})$. The proportionality constant is usually denoted by $\zeta$ and is called the impedance of the boundary.
Sound soft boundary conditions model obstacles that fully absorb the amplitude of the incoming wave, such as anechoic chambers. Sound hard boundary conditions model obstacles that fully reflect the incoming wave, such as rigid walls. Impedance boundary conditions model obstacles that partially absorb and partially reflect the incoming wave, such as porous materials.
@@ II Nematic Helmholtz–Korteweg equation
where $c_{1},c_{2}>0$ are constitutive constants. This approach was extended to the case of a general acoustic wave in Selinger et al. [25], where the elastic energy augmentation is assumed to couple the nematic director field with the gradient of the density field. In this context Virga’s theory regards liquid crystals as Korteweg fluids and proposes the constitutive relation (3) for the Cauchy stress tensor, which in a hyperelastic framework can be derived from the elastic energy density $W(\rho,\nabla\rho,\vec{n})$ [6]. The acoustic energy $W_{a}$ is derived from $W$ by assuming that the density $\rho$ represents a planar wave.
@@
Under this hypothesis, and the additional assumption that the degree of orientational order is constant, describing the nematic order with a director field $\vec{n}$ or a $Q$-tensor are equivalent.
@@
[…]
    ${\color[rgb]{0.2980392156862745,0.4470588235294118,0.6901960784313725}\definecolor[named]{pgfstrokecolor}{rgb}{0.2980392156862745,0.4470588235294118,0.6901960784313725}-i\omega\rho_{0}(\vec{v}\cdot\vec{\nu})}=c_{0}^{2}\partial_{\vec{\nu}}S(\vec{x})$ $-\rho_{0}^{2}u_{1}\partial_{\vec{\nu}}\Delta S(\vec{x})$ (31)
@@ III Plane waves
$1+\kappa^{2}+\frac{1}{4}\tau_{1}^{2}\omega^{2}\kappa^{4}=0,\qquad\tau_{1}\coloneqq 4\frac{\rho_{0}\sqrt{u_{1}}}{c_{0}^{2}}.$ (38)
@@
$1+\left(\frac{c_{0}}{c}\right)^{2}+\frac{1}{4}\tau_{1}^{2}\omega^{2}\left(\frac{c_{0}}{c}\right)^{4}=0.$ (44)
@@
$1-\left(\frac{c_{0}}{c}\right)^{2}\alpha+\frac{1}{4}\tau_{1}^{2}\omega^{2}\left(\frac{c_{0}}{c}\right)^{4}\alpha^{2}=0.$ (47)
@@
$\omega^{2}+c_{0}^{2}k^{2}+\rho_{0}^{2}u_{1}k^{4}+\rho_{0}^{2}u_{2}k^{4}\left(\vec{d}\cdot\vec{n}\right)^{2}=0,$ (51)
@@
$\omega^{2}+c_{0}^{2}k^{2}+\left[\rho_{0}^{2}u_{1}+\rho_{0}^{2}u_{2}\cos^{2}(\xi)\right]k^{4}=0.$ (52)
@@
$1+\kappa^{2}+\frac{1}{4}\tau_{2}^{2}\omega^{2}\kappa^{4}=0,$ (54)
@@ IV Plane wave reflection
    For sound-soft boundary conditions along the real axis we have $s_{0}e^{ikd_{1}x_{1}}+s_{0}Ae^{ikd_{1}x_{1}}=0$, thus we need to impose A=1A=-1.
@@
Figure 2: An acoustic reflection phenomenon in a nematic Korteweg fluid can be caused by a discontinuity in the nematic director field. We consider a Gaussian beam travelling upwards in a semicircular domain, with two different nematic director fields. (Left) The nematic director field is parallel to the wave-vector on the left side of the line marked in green, and orthogonal on the right side. As a result, a plane wave is partially reflected along the line. (Right) No reflection occurs when the nematic director has no discontinuity. The discontinuity in the nematic director field presented in the left figures cannot be achieved if the nematic director field is a general solution of the Euler–Lagrange equations associated with the Oseen–Frank energy density [28], and should be regarded as an approximation of a solution of the Oseen–Frank model exhibiting a sharp transition between two constant orientations.
@@ V Scattering by a circular obstacle
$c_{0}^{2}\Delta S^{+}_{\ell}(\vec{\xi})+\Delta^{2}S^{+}_{\ell}(\vec{\xi})+\gamma^{-1}\nabla\!\cdot\!\nabla\left[{\vec{n}\cdot\underline{\underline{\mathcal{H}S^{+}_{\ell}}}\vec{n}}\right]=0.$ (67)
@@
$c_{0}S^{+}_{\ell}(\vec{\xi})+\Delta S^{+}_{\ell}(\vec{\xi})+\gamma^{-1}{\vec{n}\cdot\underline{\underline{\mathcal{H}S^{+}_{\ell}}}\vec{n}}=0,$ (68)
@@
$c_{0}S^{+}_{\ell}(\vec{\xi})+\nabla\cdot\left[(I+\gamma^{-1}\vec{n}\otimes\vec{n})\nabla S^{+}_{\ell}\right]=0.$ (69)
@@
where $(r,\theta)$ are the polar coordinates of $\vec{x}$, $R$ is the radius of the circular obstacle, $H_{j}^{(1)}$ are the Hankel functions of the first kind, $k$ is the wave number, and $a_{j}$ are the coefficients obtained by expanding the incoming plane wave $S^{-}$ in circular harmonics via the Jacobi–Anger formula [30, eq. (22)], i.e.
@@
where $J_{j}$ is the Bessel function of the first kind and $\psi$ is the angle of incidence of the incoming plane wave. Notice now that the wave number $k$ appearing in (75) depends on the incoming plane wave and needs to satisfy the dispersion relation (51). In particular, as discussed in Section III, the wave number $k$ is smaller when the wave propagates orthogonal to the nematic director field. In regions where the incident plane wave is orthogonal to the nematic director field we can expect the amplitude of the reflected wave to be smaller, given the following asymptotic expansion of $S^{+}$ [30, eq. (38)]:
@@ Acknowledgements.
This work was funded by the Engineering and Physical Sciences Research Council [grant numbers EP/R029423/1 and EP/W026163/1], and by the Donatio Universitatis Carolinae Chair “Mathematical modelling of multicomponent systems”. The authors would like to thank E. Virga for helpful discussions and comments on the first draft of this manuscript. The authors would like to thank also T. van Beeck and C. Lehrenfeld for their help in understanding the correct boundary conditions for the Helmholtz–Korteweg equation. For the purpose of open access, the author has applied a CC BY public copyright licence to any author accepted manuscript arising from this submission. No new data were generated or analysed during this work.